myininaya (myininaya):

joe that thingy that we were talking about earlier i proved it by induction is there another way?

OpenStudy (anonymous):

i'll post what i turned in for my test. one sec, let me find it.

myininaya (myininaya):

im fixing to switch to my laptop since my desktop isn't portable

OpenStudy (zarkon):

Apparently I missed a nice problem?

OpenStudy (anonymous):

Product and sum of the roots of a polynomial.

OpenStudy (anonymous):

its part b) somewhere in the middle of page 2. The original problem was: Given a polynomial: \[a_nx^n+a_{n-1}x^{n-1}+\cdots +a_1x+a_0 = 0\] Prove that the sum of the roots is: \[-\frac{a_{n-1}}{a_n}\] for the curious, part a was to prove the Rational Roots Theorem,

OpenStudy (anonymous):

an inductive proof sounds really interesting...im gonna look into that later :)

myininaya (myininaya):

we want to show this (where r1,r2,..,rn are roots of Pn) \[P_n=(x-r_1)(x-r_2)(x-r_3) \cdots (x-r_n)\] can be wriiten as \[P_n=x^n-(r_1+r_2+r_3+ \cdots +r_n)x^{n-1}+b_{n-2}x^{n-2}-b_{n-3}x^{n-3}+ \cdots +(-1)^n(r_1 r_2 r_3 \cdots r_n)\] where the b's are some coefficients that we dont really care about \[P_1=x-r_1\] has root r1 \[P_2=(x-r_1)(x-r_2)=x^2-(r_1+r_2)x+r_1 r_2\] anyways lets assume it is true for k for even k first then we have \[P_k=(x-r_1)(x-r_2) \cdots (x-r_k)=x^n-(r_1+r_2+r_3+ \cdots r_k)x^{k-1}+b_{k-2}x^{k-2} \cdots +(r_1 r_2 \cdots r_k)\] now we want to show it is true for k+1 where k is even \[P_{k+1}=(x^k-(r_1+r_2+\cdots +r_k)x^{k-1}+gibberish+ \cdots + (r_1 r_2 \cdots r_k))(x-r_{k+1})\] \[=x^{k+1}-(r_1+r_2+ \cdots + r_k)x^k+(gibberish)x+ \cdots (r_1 r_2 \cdots r_k)x\]\[-r_{k+1}x^{k}+r_{k+1}(r_1+r_2+ \cdots + r_k)x^{k-1}+(gibberish)r_{k+1}+ \cdots -(r_1 r_2 \cdots r_k r_{k+1})\] now all we have to do is put it tgether

myininaya (myininaya):

so we can do this for odd k too

OpenStudy (anonymous):

right right. nice :)

OpenStudy (anonymous):

At first i had some points taken off my proof because I didnt show that a polynomial with roots r_1......r_n has a unique representation.

myininaya (myininaya):

thats madness lol

OpenStudy (anonymous):

i ended up proving it to him in his office when we were talking about the test so he gave them back <.< lolol I wouldnt have run into that problem if i had done it this way.

myininaya (myininaya):

this way isn't long if you are writing it down on paper it took me forever to use that equation thing

OpenStudy (anonymous):

i was waiting in suspense!! lol

myininaya (myininaya):

and after all that i couldn't brink my else to put my like terms togeher but you guys know how to do that

myininaya (myininaya):

bring*

OpenStudy (anonymous):

very nice indeed, i'll have to work it out myself so i dont forget it.

myininaya (myininaya):

induction when possible to use i think is the easiest method

OpenStudy (anonymous):

I know of a problem thats not better with induction >.< i found that out the hard way lolol

myininaya (myininaya):

show me

OpenStudy (anonymous):

prove that: \[\left(\begin{matrix}n \\ 0\end{matrix}\right)+\left(\begin{matrix}n \\ 1\end{matrix}\right)+\left(\begin{matrix}n \\ 2\end{matrix}\right)+\cdots \left(\begin{matrix}n \\ n\end{matrix}\right) = 2^n\]

myininaya (myininaya):

oh it may not be easy to use induction here lol

myininaya (myininaya):

but i might try to try it

OpenStudy (zarkon):

binomial theorem

myininaya (myininaya):

brb

OpenStudy (anonymous):

its possible with induction using the fact that: deleted since you are going to try it. but there is a much much much much much nicer way to do it.

OpenStudy (anonymous):

which zarkon spoiled!!!

OpenStudy (zarkon):

lol...sorry

OpenStudy (anonymous):

lolol, np, the Induction proof is important to know as well. its helped me with other similar problems.

OpenStudy (zarkon):

I just wrote out the induction proof and it is really not that bad

OpenStudy (zarkon):

just need to use \[{n+1\choose k}={n\choose k-1}+{n\choose k}\]

myininaya (myininaya):

i was just trying to use \[\left(\begin{matrix}n \\ k\end{matrix}\right)=\frac{n!}{k!(n-k)!}\] i should probably just use what you wrote zarkon lol

OpenStudy (zarkon):

I think it would make life easier :)

myininaya (myininaya):

yes indeed

OpenStudy (anonymous):

the inductive proof isnt bad, but when you see the binomial theorem proof your just like ".........>.>"